### v2

Abstract
Expert parallelism has been introduced as a strategy to distribute the computational workload of sparsely-gated mixture-of-experts (MoE) models across multiple computing devices, facilitating the execution of these increasingly large-scale models. However, the All-to-All communication intrinsic to expert parallelism constitutes a significant overhead, diminishing the MoE models’ efficiency. Current optimization approaches offer some relief, yet they are constrained by the sequential interdependence of communication and computation operations. To address this limitation, we present a novel shortcut-connected MoE (ScMoE) architecture with an overlapping parallel strategy, which effectively decouples communication from its conventional sequence, allowing for a substantial overlap of 70% to 100% with computation. When compared with the prevalent top-2 MoE architecture, ScMoE demonstrates training speed improvements of 30% and 11%, and inference improvements of 40% and 15%, in our distributed environments with PCIe and NVLink hardware, respectively, where communication constitutes 60% and 15% of the total MoE time consumption. Building on the ScMoE architecture, we further implement an expert offloading strategy to facilitate memory-limited inference, optimizing latency through the overlap of expert migration. Additionally, extensive experiments and theoretical analyses indicate that ScMoE not only achieves comparable but in some instances surpasses the model quality of existing approaches.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
In recent years, Transformer-based large language models (LLMs) have significantly propelled the fields of Natural Language Processing Vaswani et al. (2017); Brown et al. (2020); Wei et al. (2022b); Ouyang et al. (2022); Wei et al. (2022a); Chowdhery et al. (2023); Achiam et al. (2023), Computer Vision Dosovitskiy et al. (2021); Liu et al. (2021), and Multimodality Lu et al. (2019); Zhou et al. (2022); Zhang et al. (2021); Zhu et al. (2023). The sparsely-gated mixture-of-experts (MoE) approach has been integral in increasing parameter counts and enhancing model performance across various modalities Shazeer et al. (2017); Riquelme et al. (2021); Mustafa et al. (2022); Jiang et al. (2024). Expert parallelism Lepikhin et al. (2021); Fedus et al. (2022) has emerged as a viable strategy to distribute MoE computations efficiently over multiple devices, synergizing with conventional parallelism techniques Hwang et al. (2022); Singh et al. (2023) such as data parallelism Rajbhandari et al. (2020; 2021) and model parallelism Narayanan et al. (2021); Smith et al. (2022).
Nevertheless, expert parallelism incurs substantial All-to-All communication overhead Lepikhin et al. (2021); Fedus et al. (2022), which can contribute to approximately 50% of the total time in intra-node multi-GPUs or multi-nodes distributed environments (see Figure 1), thus forming a critical bottleneck in scaling MoE models Nie et al. (2022); Hwang et al. (2022); Mayer & Jacobsen (2020); Smith et al. (2022). Despite existing optimizations such as hierarchical All-to-All He et al. (2022); Nie et al. (2022) and pipelining Hwang et al. (2022); Zhang et al. (2023) strategies that mitigate communication delays and partially overlap communication with computation, the communication challenge persists due to the inherent sequential dependencies between these operations Wang et al. (2023). To address this constraint, our intuitive idea is to reconstruct the inputs of MoE layer by incorporating not only the current-layer but also the preceding-layer representations through a shortcut connection, thereby refining the communication dependencies and expanding the potential for optimization.
Figure 1: The overhead of MLP and top-2/top-1 MoE in a transformer block of SwinV2-MoE-S Hwang et al. (2022) model, allocating one expert per GPU with expert parallelism. The All-to-All communication takes up 60% of total time on a single node with 8$\times$×A30 GPUs, but drops to 15% on 8$\times$×A800 due to the latter’s 6$\times$ higher bandwidth provided by GPU-to-GPU NVLink Foley & Danskin (2017). Despite benefiting from NVLink, communication still approaches 50% due to the lower-bandwidth inter-node Ethernet Li et al. (2020) when scaling across multiple nodes.
In this paper, we introduce the shortcut-connected MoE (ScMoE) architecture, which completely decouples communication processes from the sequence of conventional MoE models. Our architecture is built upon the standard top-2 MoE (see Figure 2 (a)) which typically substitutes the Multi-Layer Perceptron (MLP) module with a top-2 gating MoE module in every second Transformer block (refer to the Transformer block with MoE module as “current layer”, and the preceding one without MoE module as “preceding layer”). Diverging from the top-2 approach, our ScMoE utilizes a top-1 MoE module to process preceding-layer representations via a shortcut connection, while employing a shared expert (an MLP module) to process current-layer representations. These two processes are independently managed in parallel, with their results integrated into the final output of the current layer. Furthermore, we select the intermediate representations between the Attention and MLP modules in the preceding layer as the input for the top-1 MoE. Our experiments on various preceding-layer representations (output, intermediate, input) indicate that this choice achieves optimal accuracy and a relatively longer overlap duration.
To efficiently overlap the decoupled communication and computation within our ScMoE architecture, we implement an adaptive overlapping parallel strategy that dynamically schedules operators based on actual performance metrics. Relative to existing optimization strategies, our approach not only doubles the overlap duration compared to the pipelining Huang et al. (2019); Hwang et al. (2022) but also realizes complete overlapping of communication in scenarios where communication time does not surpass the computation duration. Furthermore, our method essentially advances the MoE architecture in algorithm aspect, which is device-agnostic to improve the efficiency of MoE model, thus ensuring a broad applicability across various hardware configurations and maintaining compatibility with current optimizations.
Additionally, existing studies Hwang et al. (2023b); Yi et al. (2023) offload expert parameters to CPU memory in memory-limited inference scenarios where GPU cannot store the full MoE model. These studies utilize information in preceding layers to predict expert selection for the current MoE layer, enabling early expert migration from CPU to GPU and overlapping it with model computation. In contrast to existing speculative expert migration methods, we implement an expert offloading strategy with overlapping determinate migration, built upon our ScMoE that inherently advances expert selection to the preceding layer.
The extensive experimental results reveal that compared with the standard top-2 MoE, our ScMoE architecture achieves training speed improvements of 30% and 11% in 8$\times$×A30-PCIe and 8$\times$×A800-NVLink scenarios characterized by high and low communication overheads, respectively, and deliver inference speed improvements of 40% and 15%. Besides, our ScMoE has been demonstrated through experiments and theoretical analysis to attain or exceed the model quality of existing methods in both vision and language downstream tasks. In memory-limited inference, our expert offloading strategy reduces peak GPU memory usage by up to 60% and decreases expert migration costs by up to 75% through overlapping with computation. In addition, we conduct an in-depth analysis of our ScMoE, investigating the proposed shortcut connection and exploring the potential reasons for its algorithmic effectiveness.
In summary, our contributions are as follows:
  • We propose the shortcut-connected MoE (ScMoE) architecture that breaks conventional interdependency between communication and computation in distributed MoE models, bypassing the restrictions imposed on current communication optimization techniques.
  • We develop an adaptive overlapping strategy for advancing expert parallelism with our shortcut-connected MoE, which significantly improves the efficiency of MoE models and ensures broad compatibility.
  • We implement an expert offloading strategy with overlapping determinate expert migration upon our ScMoE, enhancing efficiency in memory-limited inference.
  • We conduct empirical evaluation and theoretical analysis on our methods, confirming that our methods accelerate MoE models while achieving comparable or even better model quality compared to existing methods, and offer in-depth analysis and discussion on the effectiveness of the proposed shortcut connection.
2 Background & Related Work
2.1 Sparsely-Gated Mixture of Experts
The Sparsely-Gated Mixture-of-Experts Shazeer et al. (2017) (MoE) layer is composed of multiple Multi-Layer Perceptron (MLP) sub-networks, termed “experts”, and employs a trainable gating network to selectively activate a subset of these experts during each iteration. Given $N$ expert networks $\{E_{i}\}_{1}^{N}$, gating network $G$ and input representation $x$, the output of MoE module can be written:
$MoE(x)=\sum_{i=1}^{N}G(x)_{i}E_{i}(x)$ (1)
Following the prevailing approach in existing MoE research, we use the noisy top-k softmax gating network to select $k$ experts for the computation, formalized by
$G(x)=Softmax(\overline{TopK}(H(x),k))$ (2)
$\overline{TopK}(H(x),k)_{i}=\begin{cases}H(x)_{i},~{}\text{if $H(x)_{i}\in TopK(H(x))$.}\\ -\infty,~{}\text{otherwise.}\end{cases}$ (3)
$H(x)_{i}=(x\cdot W_{gate})_{i}+\epsilon_{i},$ (4)
$\epsilon_{i}=StandardNormal()\cdot Softplus((x\cdot W_{noise})_{i}),$ (5)
where $\epsilon$ is tunable Gaussian noise, $W_{gate}$ and $W_{noise}$ denote two trainable weight matrices.
Leveraging sparse output of $G(x)$, this approach significantly increases the number of model parameters without causing a proportional increase in computational demand. The value of $k$ can be set to 1 or 2 or even higher values. Opting for a larger $k$ moves the model closer to the dense architecture, which generally results in higher prediction accuracy Riquelme et al. (2021), but also leads to larger execution overhead.
Figure 2: Illustrations of the standard top-2 MoE architecture (a) and the corresponding shared-expert MoE architecture (b). “SE” in (b) denotes the shared expert.
Figure 2 (a) illustrates the prevailing top-2 MoE architecture. Each Transformer block with MoE module, denoted by a light blue block and referred to as “Block-MoE,” replaces the MLP with a set of experts (“E1, E2, E3, E4”) and a gating network (“Gate”). Following prior work Lepikhin et al. (2021); Du et al. (2022); Shen et al. (2023); Hwang et al. (2022), the “Block-MoE” is interspersed with the conventional Transformer block, depicted as a gray block and referred to as “Block-MLP.”
Shared Expert. In contrast to the standard top-2 MoE architecture, the shared-expert MoE incorporates a fixed dense MLP module to process all input tokens, combining its output with the result from the top-1 gating expert for each token, as illustrated in Figure 2 (b). Given the shared expert $SE$, the output of the MoE module is formulated as:
$MoE(x)=SE(x)+\sum_{i=1}^{N}G(x)_{i}E_{i}(x)$ (6)
This method, initially proposed by DeepSpeed-MoE Rajbhandari et al. (2022), activates the same number of experts as standard MoE for computation while reducing dynamic expert selection and communication volume. Extensive empirical results from DeepSpeed-MoE and subsequent studies Dai et al. (2024); Team (2024) demonstrate that the shared expert achieves on-par or even better model quality,leading to its increasing application Xue et al. (2024); Wu et al. (2023); Chen et al. (2024); Gou et al. (2023); Gao et al. (2024).
2.2 Expert Parallelism
Figure 3: Illustration of scaling MoE transformer layer across multiple devices with expert parallelism.
Figure 4: Illustrations of various ScMoE architectures with shortcut connections to different positions of the preceding layer: (a) “Pos-1” output, (b) “Pos-2” intermediate, and (c) “Pos-3” input. The red line indicates the transmission of the preceding-layer representations to the MoE via a shortcut connection. To streamline the explanation, details regarding pre-layer normalization (Pre-Norm) and dropout procedures have been excluded. The ScMoE (Pos-2) configuration (b) is selected as the default ScMoE architecture for further implementations and experiments due to its superior performance, as shown in Table 1.
To facilitate efficient distributed training and inference of MoE models, expert parallelism is proposed to allocate unique experts to each distributed computing device such as GPU and TPU, and map tokens to their corresponding experts through All-to-All communication across the participating devices Lepikhin et al. (2021); He et al. (2021); Nie et al. (2022). As illustrated in Figure 3, the workflow of MoE employing expert parallelism is segmented into the following sequential operations: gate routing, input encode, All-to-All dispatch, expert computation, All-to-All combine, and output decode. To enhance the efficiency, input encode is employed to aggregate the token data layout to a contiguous format before All-to-All dispatch, and output decode is the inverse process after All-to-All combine. Furthermore, the integration of expert parallelism with other parallelisms Hwang et al. (2022); Singh et al. (2023); Fedus et al. (2022); Zheng et al. (2022) has been explored to support the scaling of larger MoE models on extensive distributed systems.
However, the All-to-All communication used for token transfer has been a primary factor limiting the efficiency of distributed MoE models, as illustrated in Fig 1. To address this issue, we enhance the MoE architecture and expert parallelism with shortcut connections, significantly overlapping communication costs with computation.
2.3 Memory-Limited Inference of MoE Models
While MoE effectively enhances LLMs in terms of model quality, it faces significant deployment challenges during on-device inference due to high memory demand. A common approach is to offload expert parameters to CPU memory Shen et al. (2022) in scenarios where GPU memory is insufficient to store the entire MoE model. Since the migration of activated expert parameters from CPU to GPU, which occurs after expert selection, blocks expert computation until the transfer is complete, existing studies Hwang et al. (2023b); Yi et al. (2023); Du et al. (2024) have explored prefetching the experts. For instance, Pre-gated MoE Hwang et al. (2023b) uses information from preceding layers to predict expert selection, allowing for preloading of expert parameters into GPU memory, as shown in Figure 7. This method enables overlapping the expert migration duration with the computation of preceding modules.
However, speculative expert migrations can suffer from estimation inaccuracies, as they deviate from the original logic of pre-trained models, potentially reducing inference accuracy. In contrast, we enhance the MoE architecture with a shortcut connection, utilizing the MoE module to process preceding-layer representations. This allows us to implement an expert offloading strategy with overlapping determinate migration, maintaining the pre-trained logic.
3 Shortcut-connected MoE Designs
In the prevailing Transformer-based model, the MoE module substitutes MLP to sequentially manipulate intermediate representations Lepikhin et al. (2021); Du et al. (2022); Shen et al. (2023), impeding the efficacy of existing optimization strategies He et al. (2022); Nie et al. (2022); Hwang et al. (2022); Zhang et al. (2023) due to the limited interaction within the MoE module. To address the aforementioned limitations, we propose the shortcut-connected MoE (ScMoE) architecture (Section 3.1), which enables optimization opportunities for computation-communication overlap for expert parallelism (Section 3.2) and determinate expert migration for memory-limited inference (Section 3.3).
Table 1: Test top-1 accuracy and overlap duration of pre-trained SwinV2-MoE-S models with various ScMoE configurations (illustrated in Figure 4). $\mathcal{T}_{Atten}$, $\mathcal{T}_{SE}$, and $\mathcal{T}_{MLP}$ represent the durations of Attention, shared expert, and MLP, respectively.
ScMoE ImageNet-1K (Acc@1$\uparrow$) Overlap Duration
Pos-1 79.14% $\mathcal{T}_{Atten}+\mathcal{T}_{SE}$
Pos-2 79.38% $\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$
Pos-3 79.20% $2\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$
3.1 Architectural Design
In this section, we introduce the shortcut-connected MoE (ScMoE) architecture. Unlike prevailing MoE architectures, illustrated in Figure 2, which focus on processing intermediate representations within the current layer (the Transformer block containing the MoE), our ScMoE processes representations from both the current and preceding layers. Specifically, ScMoE employs a top-1 MoE module to handle representations from the preceding layer via a shortcut connection, while a shared expert processes the current-layer representations. These two operations are conducted independently, with their outcomes integrated into the final output of the current layer, facilitating communication and computation overlap between these two processes.
While the shared expert processes the same intermediate representations in the current layer as the prevailing MoE approaches, we explore three distinct preceding-layer representations for ScMoE’s top-1 MoE process, as illustrated in Figure 4. The configurations “Pos-1” (a), “Pos-2” (b), and “Pos-3” (c) represent shortcuts connecting different positions of the preceding layer: output, intermediate, and input, respectively.
Based on our experimental results presented in Table 1, the configuration “Pos-2,” which selects intermediate representations between the Attention and MLP modules in the preceding layer as input for the top-1 MoE module, achieves optimal accuracy (79.38%) and the second longest overlap duration ($\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$). Consequently, we have chosen ScMoE (Pos-2) as the default ScMoE architecture for subsequent implementations and experiments. This ScMoE architecture can be formulated as follows:
Block-MoE:
$\mathcal{H}_{l+1}^{\text{ScMoE}}$ $=\mathcal{H}^{MH}_{l+1}+SE^{(l+1)}(\mathcal{H}^{MH}_{l+1})$ (7)
$+\sum_{i=1}^{N}G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l})$
$\mathcal{H}^{MH}_{l+1}=\mathcal{H}^{MLP}_{l}+\text{MultiHead}^{(l+1)}_{MoE}(\mathcal{H}^{MLP}_{l})$ (8)
Block-MLP:
$\mathcal{H}^{MLP}_{l}=\mathcal{H}^{MH}_{l}+\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l})$ (9)
$\mathcal{H}^{MH}_{l}=\mathcal{H}_{l-1}+\text{MultiHead}^{(l)}_{MLP}(\mathcal{H}_{l-1})$ (10)
where $\mathcal{H}_{l+1}^{\text{ScMoE}}$ refers to the output from the MoE sub-layer, $\mathcal{H}_{l+1}^{MH}$ signifies the output from the Multi-Head Attention (MultiHead) sub-layer $\text{MultiHead}^{(l+1)}_{MoE}(\cdot)$ in the $(l+1)$-th Transformer block (“Block-MoE”). $SE^{(l+1)}(\cdot)$ denotes the shared expert while $E_{1},...,E_{n}$ represent the $N$ gate-selected experts. The gating network $G(\cdot)$ is referred to as Equation 2. $\mathcal{H}_{l}^{MLP}$ and $\mathcal{H}_{l}^{MH}$ are the outputs of the MLP sub-layer $\text{MLP}^{(l)}(\cdot)$ and the MultiHead sub-layer $\text{MultiHead}^{(l)}_{MLP}(\cdot)$, respectively, in the $l$-th Transformer block (“Block-MLP”). Note that we omit the pre-layer normalization and dropout for simplicity.
In addition, we provide an in-depth theoretical analysis of our proposed ScMoE architecture in Appendix A.1, elucidating the propagation of gradients to guarantee consistent training and preserve model quality.
3.2 Overlapping Strategy for Expert Parallelism
As mentioned in the previous section, the MoE operations in ScMoE architecture are completely decoupled from the backbone network, enabling parallel execution across two independent streams: one for the shared expert process and another for the MoE process. To enhance the efficiency, we implement asynchronous All-to-All communication operators to enable the overlapping of communication and computation within these streams, while computation operators are unable to execute concurrently due to the constraints on computing resources.
Figure 5: An overview of advanced expert parallelism using our ScMoE architecture and overlapping strategy. The red line represents the decoupled MoE stream and the numbers ① through ④ denote the potential locations for expert computation.
Figure 6: The timeline of different MoE architectures with corresponding parallel strategies, including pipeline and our proposed Overlapping. In each timeline, the length of each operator represents its time cost, while multiple rows indicate the utilization of parallel CUDA streams. The standard MoE utilizes top-2 gating, whereas the shared-expert MoE and ScMoE activate one shared expert alongside one gate-selected expert.
Adaptive Operators Scheduling. We observe that operator execution times are influenced by the specific model and hardware configuration, necessitating the implementation of adaptive scheduling for operators.
Following the execution order in the MoE stream, we can directly schedule the gate routing and encode operators at the earliest viable position while deferring the decode operator to the latest position, thereby maximizing the potential duration for overlapping. Then, this challenge is distilled into the selection of an optimal position for expert computation among four possible locations ①②③④ within the shared expert stream, as depicted in Figure 5.
Formally, we define the communication costs associated with “All-to-All Dispatch” and “All-to-All Combine” as $\mathcal{T}_{disp}$ and $\mathcal{T}_{comb}$, respectively. The variable $\mathcal{K}$ is designated to represent the specific location where expert computation is applied. Prior to the expert computation, the computational costs are denoted as $\mathcal{T}_{comp}^{pre}\coloneqq\{COMP_{1},...,COMP_{\mathcal{K}-1}\}$, while the costs following the expert computation are represented as $\mathcal{T}_{comp}^{post}\coloneqq\{COMP_{\mathcal{K}+1},...,COMP_{4}\}$. Consequently, the minimal aggregate time cost for each pair consisting of one Block-MLP and one Block-MoE is
$\mathcal{T}_{overall}^{block}$ $=\min_{\mathcal{K}}(|\mathcal{T}_{comp}^{pre}-\mathcal{T}_{disp}|+|\mathcal{T}_{comp}^{post}-\mathcal{T}_{comb}|)$ (11)
$=\min_{\mathcal{K}}(|\sum_{i=1}^{\mathcal{K}-1}$ $COMP_{i}-\mathcal{T}_{disp}|+|\!\!\sum_{i=\mathcal{K}+1}^{4}\!\!\!COMP_{i}-\mathcal{T}_{comb}|)$
$\mathcal{T}_{overall}^{block}\geq|(\mathcal{T}_{comp}^{pre}+\mathcal{T}_{comp}^{post})-(\mathcal{T}_{disp}+\mathcal{T}_{comb})|$ (12)
$\mathcal{T}_{overall}^{block}\leq(\mathcal{T}_{comp}^{pre}+\mathcal{T}_{comp}^{post})+(\mathcal{T}_{disp}+\mathcal{T}_{comb})$ (13)
To demonstrate the efficiency, we have illustrated the operational timelines of various MoE architectures alongside their respective parallel strategies in Figure 6, exemplified by the selection of location ② for expert computation. Each timeline’s operator length corresponds to its execution time, and the presence of multiple rows signifies the utilization of parallel CUDA streams.
The widely-used pipeline parallel strategy equally segments input tokens into smaller fine-grained chunks, enabling concurrent computation and communication dispatched on distinct GPU streams Hwang et al. (2022); Zhang et al. (2023). Contrary to standard MoE with pipelining (2nd timeline), our ScMoE with our proposed overlapping strategy (4th timeline) reduces the total communication time by half, leading to the same cost as the shared-expert MoE (3rd timeline). Our approach further optimizes performance by overlapping communication with the computation duration ($\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$), which extends beyond the duration of expert computation achieved through pipelining.
Our strategy possesses the capability to fully overlap communication if the communication tasks can be accommodated within the overlapping window. This advantage is not shared by the pipeline strategy as it cannot overlap initial and terminal data transmissions Huang et al. (2019); Narayanan et al. (2019). In cases where communication durations exceed the available overlap duration, our strategy can be augmented with pipelining (5th timeline), thereby utilizing the expert computation duration to further hide communication.
3.3 Expert Offloading Strategy for Memory-Limited Inference
Figure 7: Illustrations of various expert migration methods to improve the efficiency of expert offloading: (a) speculative expert migration, exemplified by Pre-gated MoE Hwang et al. (2023a), and (b) our ScMoE’s determinate expert migration. The red dashed line indicates expert selection and the transfer of expert parameters from CPU memory to GPU memory, while the black or red solid lines represent the data flow of representations processed by the Attention, MLP, and expert modules.
As discussed in Section 2.3, MoE models encounter significant deployment challenges during on-device inference due to their model size exceeding GPU memory capacity, necessitating the offloading of expert parameters to CPU memory. Moreover, decoder-only models use an autoregressive process for natural language generation (NLG) inference tasks, allowing for per-token processing of MoE. Specifically, only the two activated experts (top-2 gating) for each token need to be transferred from CPU to GPU memory for computation, thereby reducing peak GPU memory usage.
To minimize the overhead caused by migrating activated expert parameters from CPU to GPU, speculative expert migration methods, represented by Pre-gated MoE Hwang et al. (2023a), use preceding-layer representations to predict expert selection, as illustrated in Figure 7 (a). This approach allows the overlap of expert migration duration with the computation of preceding modules. Moreover, speculative expert migration methods adjust only the expert selection process, while expert computation continues along the same data flow of representations as in standard MoE.
In contrast, our proposed ScMoE architecture utilizes the gate-selected expert to compute the preceding-layer representations, inherently facilitating early expert migration well before the expert computation in the current layer. Based on this, we implement an expert offloading strategy that keeps non-expert and shared expert modules in GPU memory while offloading other gate-selected experts to CPU memory. After the Attention module in the preceding layer generates intermediate representations, the gate determines expert selection and issues asynchronous migration of the activated expert, as illustrated in Figure 7 (b). This approach allows expert migration to overlap with the computation duration ($\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$). Importantly, expert selection in our method adheres to the logic of the pre-trained ScMoE model, without any speculation.
Additionally, existing expert migration methods cannot be adapted to overlap communication in expert parallelism. This is because they do not decouple dependencies in the data flow of expert processing representations, and therefore cannot adjust the All-to-All communication of these representations.
4 Experiments
4.1 Experimental Setup
Hardware Configurations. To assess the effectiveness of our proposed overlapping strategy for enhancing expert parallelism, we conduct experiments on three hardware configurations: 8$\times$×A30-PCIe, 8$\times$×A800-NVLink and 16$\times$×A800-NVLink (across 2 nodes). These configurations cover scenarios with both high and low communication-to-computation ratios. Additionally, we evaluate our proposed expert offloading strategy on a configuration with a single A30-PCIe GPU.
Experiments on Vision Model. To evaluate the efficacy of our MoE architectures on vision tasks, we conduct experiments on SwinV2-MoE model, which is a state-of-the-art vision transformer model built upon the Tutel MoE framework Hwang et al. (2022); Liu et al. (2021). Specifically, we pre-train the SwinV2-MoE models with various MoE architectures on ImageNet-1K image classification dataset, and subsequently evaluate their accuracy on the corresponding test set. It is noteworthy that the integration of the MoE module within SwinV2 is confined to stages 3 and 4, with our architectural enhancements being selectively applied to the MoE modules in stage 3—the deepest submodel. Given our hardware constraints, we configure each MoE module with 8 experts, assigning one expert per GPU device.
Experiments on Language Model. For natural language generation (NLG) tasks, we utilize the standard implementations of GPT-2 Radford et al. (2019) and GPT-3 Brown et al. (2020) from Fairseq Ott et al. (2019), augmented with Tutel MoE to construct GPT2-MoE and GPT3-MoE models featuring an 8-expert MoE module. Specifically, we implement it by substituting the MLP with MoE in the second transformer block of every consecutive pair. Then, we pre-train the models with different architectures on the OpenWebtext dataset Gokaslan & Cohen (2019). In addition to assessing the pre-training validation perplexity, we also conduct a zero-shot evaluation of the pre-trained model on the WikiText-103 Merity et al. (2016) dataset.
More setup details are illustrated in Appendix A.6
Table 2: Test top-1 accuracy and end-to-end speedup of train and inference (one iteration) for SwinV2-MoE-S Hwang et al. (2022) models with various architectures pre-trained on ImageNet-1K for 90 epochs in 8$\times$×A30-PCIe scenario, using standard MoE with top-2 gating as the baseline.
Model ImageNet-1K Train Inference
[…]
Standard top-2 MoE 79.33% 1 1
Standard top-1 MoE 78.95% 1.27$\times$ 1.39$\times$
Shared-Expert MoE 79.53% 1.24$\times$ 1.35$\times$
Our ScMoE 79.38% 1.43$\times$ 1.66$\times$
Table 3: Comparison of zero-shot perplexity on WikiText-103 and end-to-end speedup analysis of train and inference (one iteration) for our pre-trained GPT2-MoE-Medium Radford et al. (2019) models with various architectures in 8$\times$×A800-NVLink scenario, using standard top-2 MoE as the baseline.
Model WikiText-103 Train Inference
[…]
Standard top-2 MoE 19.18 1 1
Shared-Expert MoE 17.94 1.04$\times$ 1.06$\times$
Our ScMoE 17.62 1.12$\times$ 1.17$\times$
4.2 Analysis of Model Quality and Efficiency in Distributed Scenarios
In this section, we assess the model quality of our proposed ScMoE, which is a novel architectural design distinct from existing MoE models. Furthermore, we evaluate the efficiency of our ScMoE models in distributed scenarios, which are accelerated through our proposed overlapping strategy for enhancing expert parallelism. To maintain the same computational volume as the standard top-2 MoE, both the experimental Shared-Expert MoE and our ScMoE utilize one shared expert and one gate-selected expert.
Refer to caption (a) 1 Node (8$\times$×A30-PCIe) Refer to caption (b) 1 Node (8$\times$×A800-NVLink) Refer to caption (c) 2 Nodes (16$\times$×A800-NVLink)
Figure 8: Overhead analysis for each pair of Block-MLP and Block-MoE within SwinV2-MoE-S model, deployed across three different distributed scenarios. “Topk” denotes the standard top-k MoE, while the one followed by the suffix “P” indicates using pipeline optimization as implemented by Tutel Hwang et al. (2022). “Top1+SE1” refers to the shared-expert MoE.
Refer to caption (a) GPT2-MoE-Small Refer to caption (b) GPT2-MoE-Medium
Figure 9: Token-wise validation perplexity curves for training MoE models with different MoE architectures.
4.2.1 Vision Model
Table 2 shows that our ScMoE and the standard top-2 MoE attain a comparable accuracy of 79.3%, while the shared-expert MoE delivers the highest accuracy, with a marginal increase of 0.2%. These methods, which utilize two activated experts, consistently outperform the standard top-1 MoE in terms of model quality, as the top-1 approach involves fewer parameters for computation.
Notably, in the 8$\times$×A30-PCIe scenario where communication overhead accounts for 60% of the total MoE time, our ScMoE model exhibits 30% speed improvement in training and 40% in inference compared to the standard top-2 MoE. Furthermore, it even surpasses the top-1 MoE by 13% in training speed and 20% in inference speed.
Consequently, our proposed ScMoE architecture improves the MoE models’ operational efficiency without compromising on model quality. Additional experiments with the SwinV2-MoE-B model demonstrate similar results, as detailed in Appendix A.4.
4.2.2 Language Model
As illustrated in Figure 9, both our ScMoE and the shared-expert MoE, which utilize one shared expert, demonstrate lower validation perplexity curves compared to the standard top-2 MoE during training on both the GPT2-MoE-Small and GPT2-MoE-Medium models. Consistently, in zero-shot evaluations on the WikiText-103 dataset (see Table 3), these two methods yield lower perplexity scores, with our ScMoE achieving 17.62 and the shared-expert MoE reaching 17.94, in contrast to the standard top-2 MoE, which results in a perplexity of 19.18.
As exemplified by the performance of GPT2-MoE-Small and GPT2-MoE-Medium models, our ScMoE performs high model quality across a range of network depths. Notably, our ScMoE demonstrates a superior training curve and zero-shot perplexity compared to the shared-expert MoE in the GPT2-MoE-Medium model, while the shared-expert MoE exhibits a slightly better training curve in the GPT2-MoE-Small model. This suggests that our ScMoE may offer greater benefits in deeper networks, aligning with our theoretical analysis of the shortcut connections, as detailed in Appendix A.1. Additional zero-shot results of GPT2-MoE-Small are presented in Appendix A.4.
In addition, our ScMoE achieves 11% speed improvement in training and 15% in inference compared to standard top-2 MoE in the 8$\times$×A800-NVLink scenario, where communication constitutes 15% of the total MoE time.
4.2.3 Analysis of Overhead and Acceleration
In addition to exhibiting the significant end-to-end speedup of our ScMoE in Tables 2 and 3, we delve into a detailed analysis of the overhead and the acceleration effect with our overlapping strategy, which can be generalized to other Transformer-based MoE models.
In the communication-intensive 8$\times$×A30-PCIe scenario (Figure 8a), our ScMoE overlaps 70% communication time, resulting in a 27% speed improvement over shared-expert MoE, a 42% improvement over the pipelined standard top-2 MoE, and a 15% improvement over the pipelined standard top-1 MoE. In the 8$\times$×A800-NVLink scenario (Figure 8b), which features almost minimal communication overhead, our approach maintains its acceleration by completely overlapping communication.
In the multi-node distributed scenario (Figure 8c), with 16$\times$×A800-NVLink across two nodes, communication incurs more significant overhead than in the single-node 8$\times$×A800-NVLink scenario due to the lower-bandwidth inter-node Ethernet Li et al. (2020). Here, our ScMoE achieves complete overlap, resulting in a 24% speed improvement over the shared-expert MoE, a 43% improvement over the pipelined standard top-2 MoE, and a 13% improvement over the pipelined standard top-1 MoE.
In general, our ScMoE delivers a significant acceleration over the standard top-2 MoE, and even outperforms the top-1 MoE when communication exceeds approximately 20% of the total MoE time. Additionally, our ScMoE can fully overlap communication in scenarios where communication does not exceed an estimated 50% of the total MoE time.
4.2.4 Analysis of More Activated Experts
As increasing the number of activated experts within standard MoE is correlated with enhancements in model quality, we implement this augmentation in our ScMoE by increasing the count of gate-selected experts that process the preceding-layer representations, while maintaining the process of current-layer representations. To investigate the benefits of more activated experts, we implement the ScMoE-2, which employs top-2 experts for the preceding layer and one shared expert for the current layer.
Comparative analyses with the standard top-3 MoE, which has the same computational volumes as our ScMoE-2, reveal that our ScMoE architectures maintain superiority in both model quality and efficiency, as evidenced in Table 4. Furthermore, akin to the standard MoE, our ScMoE consistently improves with additional expert activation, shown by a decrease in validation perplexity from 16.46 with ScMoE to 16.27 with ScMoE-2.
Although activating more experts incurs higher time costs, the efficiency improvements of our overlapping strategy remain significant. For instance, our ScMoE-2 requires merely 95% and 93% of the time cost necessary for the standard top-2 MoE respectively in training and inference, despite processing increased computational loads.
Table 4: Comparison of validation perplexity and end-to-end speedup analysis of train and inference (one iteration) for our pre-trained GPT3-MoE-XL Brown et al. (2020) models with various architectures in 8$\times$×A800-NVLink scenario, using standard MoE with top-2 gating as the baseline. “ScMoE-2” refers to the activation of one shared expert and two gate-selected experts.
Model Validation Train Inference
[…]
Standard top-2 17.52 1 1
Our ScMoE 16.46 1.12$\times$ 1.18$\times$
[…]
Our ScMoE-2 16.27 1.05$\times$ 1.08$\times$
4.3 Analysis of Memory-Limited Inference
To tackle the challenges of on-device inference, where it is not feasible to store the entire MoE model parameters in GPU memory, our strategy involves offloading gate-selected experts to CPU memory while retaining non-expert and shared expert modules in GPU memory. As demonstrated in Table 10 (a), our expert offloading strategy reduces peak GPU memory usage by 50% for the GPT2-MoE-Medium model and by 60% for the GPT3-MoE-XL model when deployed in the inference scenario using a single A30-PCIe GPU. Furthermore, it is anticipated that models with more gate-selected experts in each MoE module will experience a larger percentage reduction in GPU memory usage.
Since the offloaded expert parameters must be loaded into the GPU memory for expert computation, the blocking execution of this expert migration results in significant overhead. As shown in Table 10 (b), the blocking expert migration introduces an additional overhead of 80% in GPT2-MoE-Medium and 240% in GPT3-MoE-XL, substantially increasing the MoE block latency. To mitigate this issue, our strategy of asynchronously executing the determinate expert migration effectively reduces the additional costs by 75% in GPT2-MoE-Medium and 25% in GPT3-MoE-XL.
Furthermore, it is evident that expanding the model size from Medium to XL significantly raises the cost proportion related to expert migration. This is because the per-token decoding process during inference is memory-bound Patel et al. (2024); Wu et al. (2024). The larger model size leads to a proportional increase in the duration of memory transfer, without a corresponding increase in computation time.
Refer to caption (a) Peak GPU Memory Usage Refer to caption (b) MoE Block Latency
Figure 10: Peak GPU memory usage (a) and MoE block latency (b) for various memory-limited inference methods applied to the GPT2-MoE-Medium and GPT3-MoE-XL models using ScMoE. “GPU-only” indicates that the entire model is stored in GPU memory. “Offload” refers to our strategy of offloading expert parameters to CPU with blocking expert migration. “Offload-Async” denotes the use of asynchronous expert migration to overlap its duration.
Refer to caption (a) Repeated selection Refer to caption (b) L2-Distance Refer to caption (c) Preceding-layer score Refer to caption (d) Current-layer score
Figure 11: Results from the analysis of the proposed shortcut connection, during the 90-epoch training (including a 10-epoch warm-up) of the SwinV2-MoE-S model Hwang et al. (2022). Employing the same MoE module to select the top-1 expert twice for processing each input token’s two representations from the current and preceding layers, respectively, (a) illustrates the percentage of tokens that retain the same expert selection across the current layer and preceding layer, (b) shows the L2 distance between these two representations. Using the DGMoE, which imposes a constraint against repeatedly selecting the same expert, (c) presents the average gating score for the preceding-layer representations, (d) displays the average gating score for the current-layer representations.
4.4 Analysis of the Proposed Shortcut Connection
Our comprehensive empirical results have demonstrated that our ScMoE architecture facilitates efficiency optimizations without compromising model quality. We further explore the proposed shortcut connection in greater depth, uncovering the reasons behind its algorithmic effectiveness.
Firstly, we investigate the use of the same MoE module to select the top-1 expert twice for processing each input token’s current-layer and preceding-layer representations, respectively. As illustrated in Figure 11 (a), we observe that the same gating network typically selects the same expert for the two representations of most tokens. As the training progresses, the token percentage of repeating selection initially escalates, peaking at 98%, and then diminishes, with a significant drop manifested in the last MoE sub-layer.
Next, we measure the L2 distance (similarity) between each token’s preceding-layer and current-layer representations. Figure 11 (b) illustrates that, as training advances, the L2 similarity initially decreases with network depth, then increases, and ultimately reaches its maximum value in the final layer. Since the gating network is used to classify the representations, this similarity may lead to the repeated selection of the same experts, as evidenced by the correlation between the results in Figure 11 (a) and (b).
Notably, we observe this experimental architecture only achieves model quality on par with the standard top-1 MoE, despite incurring the same computational cost as the top-2 MoE. To address this, we impose a constraint on the MoE module to ensure that it selects a different expert for the current layer than the one chosen for the preceding layer. We refer to this enhanced experimental architecture as DoubleGating MoE (DGMoE), which achieves model quality comparable to standard top-2 MoE by imposing the constraint against selecting the same expert repeatedly.
With this constraint, we observe gating score behaviors are similar to those of the standard top-2 MoE Riquelme et al. (2021), as illustrated in Figure 11 (c) and (d). Specifically, in (c), the gating scores associated with representations from preceding layers grow quickly initially, converging thereafter, and achieving their maximum in the final MoE sub-layer. By comparison, in (d), the gating scores of current layer across all MoE sub-layers are around $0.2$ (±$\pm 0.04$), lower than the $0.2$ to $0.8$ range for that of preceding layer.
In summary, the experiments indicate that the combined results of two representations, produced by the same expert, are approximately equivalent to those obtained by using a single expert to process only the current-layer representations. This equivalence may arise from the similarity between the preceding-layer and current-layer representations of each token. Furthermore, enforcing two distinct experts can enable the DGMoE architecture to achieve behaviors and model quality similar to those of a standard top-2 MoE.
Analogously, our ScMoE architecture has already allocated distinct experts for the two representations of each token, thereby maintaining a similar behavior to the shared-expert MoE architecture and ensuring comparable model quality. Given the differences between standard MoE and shared-expert MoE, we have observed that DGMoE achieves accuracy levels closer to those of the standard top-2 MoE, while ScMoE aligns more closely with shared-expert MoE, as demonstrated in the results presented in Appendix A.2.
Additionally, our ScMoE model may leverage the differences between the representations of preceding layers and those of the current layer, while still preserving significant similarities. This distinction potentially provides additional information that enhances the training of MoE models, especially considering that MoE models typically require larger training datasets compared to dense MLP models Artetxe et al. (2022); Fedus et al. (2022); Xue et al. (2022).
5 Conclusion
The inherent dependencies of communication in distributed MoE models hinder parallel optimization techniques to improve execution efficiency. To overcome this, we propose a novel shortcut-connected MoE (ScMoE) architecture, and develop a communication overlapping parallel strategy. Through comprehensive empirical evaluation and theoretical analysis, our approaches demonstrate better execution efficiency while maintaining, or exceeding, the model quality of existing methods. Moreover, we further implement an expert offloading strategy upon our ScMoE to enable memory-limited inference and optimize latency by overlapping expert migration. Additionally, we provide insightful analysis and discussion on our proposed MoE architecture.
Appendix A Appendix
A.1 Theoretical Analysis
In this section, we delve deeper into the understanding of our proposed shortcut-connected MoE (ScMoE) architecture, presenting a theoretical foundation focused on the propagation of gradients to guarantee consistent training and preserve model quality. Our analysis is confined to our ScMoE architecture as depicted in Figure 4(b); however, the same principles and derivations can be easily extended to other shortcut-connected MoE architectures. Building upon Equations 7 to 10, we can derive
$\mathcal{H}_{l+1}$ $=\mathcal{H}^{MH}_{l}+\Big{(}\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l})$ (14)
$+\text{MultiHead}^{(l+1)}(\mathcal{H}^{MH}_{l}+\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l}))$
$+\text{SE}^{(l+1)}(\mathcal{H}^{MH}_{l}+\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l})$
$+\text{MultiHead}^{(l+1)}(\mathcal{H}^{MH}_{l}+\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l})))$
$+\sum_{i=1}^{N}G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l})\Big{)}$
$\mathcal{H}^{MH}_{l}$ $=\mathcal{H}_{l-1}+\text{MultiHead}^{(l)}(\mathcal{H}_{l-1})$ (15)
It is observable that Equations 14 and 15 share an identical structural expression. Consequently, we consider each pair of Block-MoE and Block-MLP layers as a single entity, and every sub-layer, denoted as $\mathcal{F}$, with its corresponding parameters $\mathcal{W}_{l}$, conforms to the equation
$x_{l+1}=x_{l}+\mathcal{F}_{\mathcal{W}_{l}}(x_{l})$ (16)
Here, $x_{l}$ represents the input, and $x_{l+1}$ represents the output of the $l$-th sub-layer. By applying this relationship recursively, the output of the uppermost $L$-th sub-layer, $x_{L}$, can be deduced as follows
$x_{L}=x_{l}+\sum_{i=l}^{L-1}\mathcal{F}_{\mathcal{W}_{l}}(x_{l})$ (17)
Let’s consider the loss function as $\mathcal{E}$. Using the chain rule, we can calculate the derivative of the loss with respect to $x_{l}$, and we have
$\frac{\partial{\mathcal{E}}}{\partial{x_{l}}}$ $=\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}\frac{\partial{x_{L}}}{\partial{x_{l}}}=\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}\Big{(}1+\frac{\partial{}}{\partial{x_{l}}}\sum_{i=1}^{L-1}\mathcal{F}_{\mathcal{W}_{i}}(x_{i})\Big{)}$ (18)
It’s clear that the additive component of the error gradient $\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}$ ensures direct information propagation back to any sub-layer $x_{l}$. Additionally, its advantage is that the number of product elements on the right side is independent of the network’s depth. Therefore, as $L$ increases, it is less likely to encounter the gradient vanishing or exploding problem, ensuring stable training and sustained performance levels in our proposed MoE architectures.
Figure 12: Illustration of the experimental DoubleGating MoE (DGMoE) architecture.
A.2 Analysis of the DoubleGating MoE (DGMoE)
To delve deeper into our architecture with shortcut connection, we introduce the DoubleGating MoE (DGMoE) architecture, which employs dual top-1 gating mechanisms to independently process the representations from the preceding and current layers, as illustrated in Figure 12. Building upon Equations 7 to 10, and contrasting with ScMoE, DGMoE can be formulated as
$\mathcal{H}_{l+1}^{\text{DGMoE}}$ $=\mathcal{H}^{MH}_{l+1}+\sum_{i=1}^{N}(G(\mathcal{H}^{MH}_{l+1})_{i}E_{i}(\mathcal{H}^{MH}_{l+1})$ (19)
$+G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l}))$
where $\mathcal{H}_{l+1}^{\text{DGMoE}}$ refers to the output from the MoE module.
However, as delineated in Equation 19, a potential issue arises when a token at the current layer selects the same top-1 expert as the preceding layer, inadvertently collapsing the intended top-2 gating mechanism into a de facto top-1 gating mechanism. To mitigate this, we introduce a constraint that ensures the activation of two distinct experts. In practice, this is achieved by first documenting the indices of experts triggered by the preceding-layer representations. Subsequently, if the preceding-layer representation coincidentally targets the same expert as the current layer, that is, if $\overline{TopK}(H(\mathcal{H}^{MH}_{l}),1)=\overline{TopK}(H(\mathcal{H}^{MH}_{l+1}),1)$, we activate the second-highest-ranking expert from the top-2 selection for the current layer, i.e., $\overline{TopK}(H(\mathcal{H}^{MH}_{l+1}),2)_{2}$.
As illustrated in Table 6 and Table 7, our DGMoE achieves comparable accuracy to the standard top-2 MoE across both vision and language tasks. Meanwhile, our ScMoE demonstrates performance more akin to the shared-expert MoE.
A.3 Analysis of Shared Expert Gate
In contrast to the gate-selected expert, the shared expert is fixed to process all representations without computing a gating score through the gating network. Therefore, common practices for shared-expert MoE employ a shared-expert gate to generate the coefficient for the shared expert’s output Team (2024); Dai et al. (2024). Specifically, the shared-expert gate is a linear layer that uses the input representation of the MoE module as its input to generate a coefficient similar to a gating score. Building upon Equation 6, the output of the shared expert can be expressed as
$SE(x)=SEGate(x)\cdot\text{MLP}(x)$ (20)
where $SEGate(\cdot)$ denotes the shared-expert gate and $\text{MLP}(\cdot)$ represents the essence of the shared expert.
As demonstrated in Table 5, the incorporation of the shared-expert gate significantly enhances model performance in our experimental vision tasks. In the absence of the shared-expert gate, the quality of MoE architectures with shared experts declines from that of a standard top-2 MoE to that of a standard top-1 MoE, despite maintaining the same computational volume as the standard top-2 MoE.
Table 5: Comparison of top-1 accuracy on the ImageNet-1K test set for SwinV2-MoE-S models, with and without the shared-expert gate (denoted as “SE-Gate”).
Model With SE-Gate Without SE-Gate
Our ScMoE (Pos-1) 79.14% 78.78%
Our ScMoE (Pos-2) 79.38% 78.98%
Our ScMoE (Pos-3) 79.20% 78.29%
Shared-Expert MoE 79.53% 79.02%
Table 6: Comparison of top-1 accuracy on the ImageNet-1K test set for SwinV2-MoE-S and SwinV2-MoE-B models with various architectures: top-2/top-1 gating standard MoE, shared-expert MoE, our DGMoE, and ScMoE, each pre-trained for 90 epochs on the ImageNet-1K classification dataset.
Model SwinV2-MoE-S SwinV2-MoE-B
(Acc@1$\uparrow$) (Acc@1$\uparrow$)
Standard top-2 MoE 79.33% 80.48%
Standard top-1 MoE 78.95% 80.05%
Shared-Expert MoE 79.53% 80.62%
[…]
Our ScMoE 79.38% 80.56%
Table 7: Comparison of zero-shot perplexity on WikiText-103 for our pre-trained GPT2-MoE-Small and GPT2-MoE-Medium models with various architectures.
Model GPT2-MoE-Small GPT2-MoE-Medium
[…]
Standard top-2 MoE 31.60 19.18
Shared-Expert MoE 29.15 17.94
[…]
Our ScMoE 29.10 17.62
A.4 Evaluation Across Different Model Sizes
Table 6 and Table 7 illustrate that our experimental MoE architectures consistently achieve analogous model quality across different model sizes, as expounded in the detailed analysis within the main body of this paper.
A.5 Share MoE Across Multiple Layers via Shortcut Connections
From a certain point of view, our shortcut-connected MoE architectures can be conceptualized as the sharing of one MoE module across multiple transformer layers. Parameter sharing across different layers has been validated as a method to enhance parameter efficiency and improve model quality, as evidenced in existing research Lan et al. (2019); Dehghani et al. (2018); Xue et al. (2022); Huang et al. (2017).
The empirical analysis of our novel MoE architectures suggests that the MoE modules shared across multiple layers via shortcuts could offer a more parameter-efficient solution. We conduct experiments on a preliminary architecture DGMoE-Share which shares a single MoE for two pairs of transformer blocks. It reduces the parameter count from 157M to 124M, while maintaining the same volume of expert computation as the standard top-1 MoE. The DGMoE-Share achieves a 78.45% accuracy on the vision task, incurring a minimal accuracy decrement of 0.5% relative to the standard top-1 MoE. We anticipate the discovery of more efficient architectures through future explorations. Additionally, the optimization of training hyperparameters for the shortcut-connected MoE requires more investigation.
Table 8: Hyperparameters for GPT-MoE models.
Parameter GPT2-MoE-Small GPT2-MoE-Medium GPT3-MoE-XL
Num. layers 12 24 24
Embedding dim 768 1024 2048
Num. attention heads 12 16 32
Num. experts per layer 8 8 8
Num. parameters 323M 1.1B 4.1B
Context/sequence length 1K 2K 2K
Training tokens 180B 180B 18B
Warmup tokens 100M 100M 10M
Batch size 256 64 32
[…]
LR-Scheduler Inverse_sqrt Inverse_sqrt Inverse_sqrt
Capacity factor 2.00 2.00 2.00
MoE loss coefficient 0.01 0.01 0.01
A.6 Experimental Details
Table 9: Hyperparameters for SwinV2-MoE models.
Parameter SwinV2-MoE-S SwinV2-MoE-B
Image size 192$\times$192 192$\times$192
Window size 12$\times$12 12$\times$12
Embedding dim 96 128
Num. layers [ 2, 2, 18, 2 ] [ 2, 2, 18, 2 ]
Num. attention heads [ 3, 6, 12, 24 ] [ 4, 8, 16, 32 ]
Num. experts per layer 8/16 8
Batch size 1024 1024
Epochs 90 90
Warmup epochs 10 10
[…]
Capacity factor 1.25 1.25
MoE loss coefficient 0.01 0.01
Table 8 summarizes the hyperparameters for training the GPT-MoE models including GPT2-MoE-Small, GPT2-MoE-Medium and GPT3-MoE-XL. We implement our models based on the Fairseq Ott et al. (2019) and Tutel Hwang et al. (2022) frameworks. And we use the Openwebtext dataset Gokaslan & Cohen (2019) for pre-training, with the GPT-2 BPE tokenizer Radford et al. (2019). In addition to evaluating the training curve, we also conduct a zero-shot evaluation on WikiText-103 Merity et al. (2016) dataset. Additionally, we maintained uniform hyperparameters across all models to ensure comparability, and we recognize that optimal parameter tuning may further enhance model performance, a subject for subsequent research.
Table 9 summarizes the hyperparameters for training the Swin-MoE models including SwinV2-MoE-S and SwinV2-MoE-B. We develop our experiments based on the models and MoE framework which are provided by Tutel Hwang et al. (2022), and also follow its hyperparameter configurations. The number of gate-selected experts per MoE module corresponds to the number of GPUs. Specifically, the experiments related to overhead and acceleration analysis in a 2-node (16×A800-NVLink) scenario utilize 16 experts per MoE module, while other cases use 8 experts. Then, we pre-train the models on ImageNet-1K image classification dataset. To maintain the comparability of our experiments, we limit our modifications solely to the MoE architectures and keep the hyperparameters and random seeds consistent. In addition, the experimental results related to efficiency are the averages of multiple samples over different periods.